OpenStudy (anonymous):

What is the solution set for |-4x - 4| - 28 = -4?

OpenStudy (anonymous):

isolate the absolute value expression by adding 28 to both sides to get: |-4x-4| = -4 + 28 |-4x-4| = 24 whenever you have an absolute value, you'll have to solve two equations. in your case it is these two: -4x-4 = 24 or -4x-4 = -24 can you solve both these equations to get your answer?

OpenStudy (anonymous):

is the answer going to be x = 7 and x = 5?